If a solution has a pH of 5, what will be its pOH?
1
Expert's answer
2020-06-04T10:37:49-0400

Solution.

"pH=5;"

"pH+pOH=14\implies pOH=14-pH;"

"pOH=14-5=9;"

Answer: "pOH=9."